For the given network, find the minimum spanning tree distance. Select one: a. 21 b. 16 c.19 d. 18
Answer #1

No of vertexes = edges - 1 = 6 - 1 = 4

1+2+4+6+5 = 18

The minimal spanning tree distance is 18